OpenStudy (anonymous):

John gave the following explanation for factoring the trinomial x2 - x - 2: "I need to find factors of -2. -1 and 2 multiply to -2 so my factors are (x - 1)(x + 2)." Part 1: Are John's factors correct? If not, explain where he made his mistake. Part 2: What could John have done to check the accuracy of his factors before submitting his Module 3 test? (4 points)

OpenStudy (ankit042):

@pmiddleton6 you know how to factor x^2-x-2?

OpenStudy (anonymous):

no please help

OpenStudy (ankit042):

you can write x^2-x -2 as x^2 -2x +x -2 x(x-2)+1(x-2)=0 (x-2)(x+1)=0 x=2 or x=-1

OpenStudy (ankit042):

For the part B you can easily check the factors by putting value of x in the equation if result is zero then factors are correct!

OpenStudy (anonymous):

okay thank you very much

OpenStudy (ankit042):

dHope you got the concept

OpenStudy (anonymous):

Could you explain part two a little bit more?

OpenStudy (ankit042):

np so f(x) = x^2 -x-2 for the values of x which are factors of the quadratic equation value of f(x) will be =0 so you can directly check by putting value of x

OpenStudy (ankit042):

In this example John got x as 1 and -2 ; so if we calculate f(1) we get 1-1-2 f(1)=-2 and f(-2)=4 so we knpw he made a mistake

OpenStudy (anonymous):

oh okay i get it!!

OpenStudy (ankit042):

Great!